### v1

# Quantifying Fire Risk Index in Chemical Industry Using Statistical Modeling Procedure

Abstract
Fire incident reports contain detailed textual narratives that capture causal factors often overlooked in structured records, while financial damage amounts provide measurable outcomes of these events. Integrating these two sources of information is essential for uncovering interpretable links between descriptive causes and their economic consequences. To this end, we develop a data-driven framework that constructs a composite Risk Index, enabling systematic quantification of how specific keywords relate to property damage amounts. This index facilitates both the identification of high-impact terms and the aggregation of risks across semantically related clusters, thereby offering a principled measure of fire-related financial risk. Using more than a decade of Korean fire investigation reports on the chemical industry classified as Special Buildings (2013–2024), we employ topic modeling and network-based embedding to estimate semantic similarities from interactions among words, and subsequently apply Lasso regression to quantify their associations with property damage amounts, thereby estimate fire risk index. This approach enables us to assess fire risk not only at the level of individual terms but also within their broader textual context, where highly interactive related words provide insights into collective patterns of hazard representation and their potential impact on expected losses. The analysis highlights several domains of risk, including hazardous chemical leakage, unsafe storage practices, equipment and facility malfunctions, and environmentally induced ignition. The results demonstrate that text-derived indices provide interpretable and practically relevant insights, bridging unstructured narratives with structured loss information and offering a basis for evidence-based fire risk assessment and management.
Keywords: Fire incident analysis; Risk index; Text mining; Chemical Special Buildings; Network interaction
1 Introduction
Fire accidents pose significant threats to human life, property, and industrial activities. In particular, facilities in the chemical industry are highly vulnerable due to the presence of flammable and hazardous materials, which can lead to catastrophic outcomes once a fire occurs. Understanding the underlying risk factors of such accidents is therefore crucial for developing effective prevention strategies, designing safety regulations, and minimizing social and economic damages.
While structured fire statistics and engineering-based indicators provide valuable information, they often fail to capture the nuanced circumstances surrounding each incident. In contrast, post-incident investigation records written by frontline fire officers contain rich textual descriptions of ignition causes, equipment failures, material involvement, and contextual human or environmental factors observed on site. These narratives represent first-hand accounts of accident circumstances, preserving details that are easily overlooked or lost when data are reduced to predefined categories. Moreover, examining textual data reveals co-occurrence patterns among words, indicating that fire-related factors often appear in dependent and interrelated structures. Such contextual associations can only be identified through systematic analysis of unstructured narratives, enabling a deeper understanding of how multiple risk factors interact to influence the likelihood and severity of fire accidents. Analyzing such textual evidence is therefore essential not only for uncovering hidden or emerging risk factors, but also for linking causal mechanisms with the resulting scale of damage.
Previous research has employed a variety of approaches to analyze fire-related risks. Text-based studies have demonstrated the potential of mining unstructured documents. For instance, Kim and Hwang [10] applied topic modeling techniques to accident verdicts from ship fire cases to identify ignition sources, flammable materials, and negligence-related causes. While this study represented an important step in applying text analytics to fire accidents, it primarily relied on word frequency clustering and did not capture the correlated structure among words that reflects the interdependent nature of fire risk factors. Similarly, Tirunagari [20] investigated maritime accident investigation reports using text mining to extract causal relations among contributing factors. Although not focused on fires, this study highlights the methodological potential of analyzing free-form investigation records, underscoring the value of unstructured narratives in uncovering complex causal structures in accident data.
Structured and indicator-based approaches have also received considerable attention. Ma et al. [14] conducted a comprehensive data-driven analysis of over one million building fire reports, integrating structured incident records with socioeconomic and structural variables to assess the effects of detection systems and extinguishing devices on fire spread and injury risk. Although similar in its goal of combining textual and structured sources, its textual component was limited to predefined categorical fields rather than free-text narratives, preventing it from capturing word co-occurrence patterns or the indexical role of terms. In a related effort, Zhang et al. [25] developed an indicator system for urban fire risk assessment, emphasizing meteorological conditions and building characteristics. Yet, this approach excluded unstructured textual data, thereby omitting rich qualitative details from fire incident reports.
Finally, research has also explored narrative-based perspectives on fire events. Russo et al. [17] analyzed wildfire narratives to identify and characterize multiple social storylines concerning causes, consequences, and potential solutions. This work illustrates how unstructured narratives can provide a richer and more contextualized understanding of fire events by linking causal descriptions to broader societal implications.
Taken together, these studies highlight the promise of both textual and structured approaches for understanding fire risks, but also reveal important limitations. Existing work has seldom analyzed fire investigation narratives written directly by frontline fire officers, which contain detailed accounts of ignition mechanisms, equipment failures, and contextual factors surrounding each incident. To address this gap, our study develops a framework that systematically analyzes such unstructured text while explicitly accounting for the dependent structure among words, thereby identifying coherent topics that represent interrelated fire causes. Furthermore, by integrating this textual analysis with structured data on fire property losses, we construct a risk index factor that quantifies the degree of fire risk across different categories of incidents. This combined approach not only facilitates the extraction of meaningful risk factors through salient keywords, but also provides researchers and practitioners with an interpretable index that reflects both causal conditions and damage outcomes. Ultimately, the proposed framework enables a more comprehensive and data-driven understanding of fire risk, helping stakeholders identify and prioritize critical factors for prevention and mitigation.
Contributions
This study makes the following contributions:
  • First large-scale analysis of Korean fire investigation narratives. To the best of our knowledge, this is the first systematic attempt to analyze over a decade (2013–2024) of textual records on the chemical industry classified as Special Buildings, written by fire officers. These narratives provide direct accounts of ignition sources and causal conditions that have not been previously examined in quantitative fire safety research.
  • Integration of textual evidence with economic loss indicators. By linking unstructured narratives with structured data on fire property damage, we move beyond frequency-based measures of risk. This integration enables the identification of risk factors that matter not only for their occurrence but also for their economic impact, thereby offering a more practical and policy-relevant assessment of fire risk.
  • Development of a statistically grounded risk index. We combine established statistical approaches [9] with a Lasso regression framework to construct a novel composite index. This allows us to capture meaningful dependency structures among words and quantify their contribution to observed loss outcomes. The resulting index provides an interpretable and efficient tool to quickly identify critical fire-related terms that elevate risk.
The remainder of this paper is organized as follows. Section 2 introduces the dataset and outlines the analytical procedures, including latent topic estimation via the Biterm Topic Model, topic clustering through a Latent Space Item Response Model, and the construction of a risk index factor using Lasso regression. Section 3 presents the results, covering topic characterization through words, thematic aggregation of topics, and evaluation of the proposed risk index factor. Section 4 discusses the implications of the findings and concludes the study.
2 Materials and Methods
Figure 1 presents the proposed analytical framework for constructing a fire risk index from unstructured fire investigation texts. In the first step (Step 1), natural language processing techniques are applied to preprocess the textual records and extract nouns, with a particular emphasis on building a domain-specific lexicon related to the chemical industry. This enables the identification and expansion of keywords that are directly relevant to chemical processes and accident scenarios, providing a structured corpus for further analysis.
In the second step (Step 2), a Biterm Topic Model [24] is employed to classify documents into latent topics and to estimate the distribution of words within each topic. This step not only offers a compact summary of large-scale documents but also transforms unstructured text into a topic–words distribution matrix that can be further exploited. Using these topic–word distributions as input for the subsequent latent space model is advantageous, as it embeds words into a representation that reflects their co-occurrence patterns, thereby facilitating the estimation of meaningful word–word interactions that would not be apparent from raw text alone.
In the third step (Step 3), the latent item response model [8] is applied to infer the positions of words in a continuous latent space, which captures their semantic relationships. The estimated distances between words can be interpreted as measures of semantic similarity: for example, if two words are placed close to each other in the latent space, they are more likely to represent semantically related concepts. Leveraging these latent positions, words are clustered into semantically coherent groups, thereby enabling the construction of interpretable clusters of risk-related vocabulary.
In the final step (Step 4), structured data on property damage amounts are incorporated into the analysis. Specifically, Lasso regression is used to estimate coefficients linking each word to the magnitude of property loss. Words with higher coefficients can thus be interpreted as risk factors associated with greater expected damages. Beyond word-level inference, this integration allows for cluster-level analysis: by examining the aggregated coefficients of words within each cluster, we can identify which semantic groupings correspond to high-risk factors in terms of potential financial losses. Taken together, this framework not only provides a systematic way to quantify fire risk from unstructured narratives but also bridges semantic information extracted from text with structured loss data to yield interpretable and practically meaningful risk indices.
Figure 1: Analytical framework for constructing a fire risk index from unstructured investigation texts. The process begins with text mining to preprocess investigation records and extract key nouns (Step 1). A Biterm Topic Model is then applied to estimate topic–word distributions (Step 2). Next, a latent item response model is used to infer topic–word interactions and to estimate latent positions of words, which are subsequently clustered to capture semantic groupings (Step 3). Finally, Lasso regression estimates the coefficients of words using property loss data as the outcome, yielding a composite risk index that integrates words’ semantic closeness with structured information derived from loss amounts (Step 4).
2.1 Data
This analysis focuses on fire accidents in Special Buildings in Korea that fall under the chemical industry (excluding plastic production) within the category of Manufacturing uses. In Korea, Special Buildings are legally designated facilities identified as high fire-risk based on their intended use, scale, and other criteria. These facilities are overseen by the Korea Fire Protection Association (KFPA) for fire prevention purposes. As of the end of 2024, there were 54,517 such buildings nationwide, and the total number changes annually.
The dataset used in this study consists of records from post-incident investigations of fire accidents in these chemical industry Special Buildings between 2013 and 2024. The fire investigation records are unstructured text documents describing the fire circumstances, such as the cause, involved equipment, and ignited materials. All investigation texts are written in Korean. In addition, the dataset includes property damage estimates provided by local fire departments, covering both movable and immovable property losses.
2.2 Latent Topic Estimation via the Biterm Topic Model
We employ the Biterm Topic Model (BTM) to uncover latent semantic structures within textual statements written by firefighters during post-incident investigations of fire causes. As a preprocessing step, morphological analysis is conducted to decompose words into their base morphemes. From these, we extract nouns in their canonical form, which constitute the initial corpus. To enrich this corpus, we expand the vocabulary set using the word2vec algorithm, which maps words into a Euclidean latent space according to semantic similarity. By identifying words in close proximity within this embedding space, we obtain an augmented corpus that better captures the semantic landscape of the texts.
The ultimate objective is to identify a collection of latent topics that summarize the semantic content of the documents. Each topic is represented as a distribution over words, while each document is modeled as a mixture of these latent topics. Since those reports are typically consist of fewer than 200 words and can thus be regarded as short texts, the BTM is particularly suitable for this setting. The model relies on extracting biterms, i.e., unordered word pairs within documents, which serve as the input to the topic model.
The BTM is founded on several key assumptions:
  • Each pair of words (biterm) is assumed to arise from an underlying latent topic.
  • Topics themselves represent semantically coherent clusters of words.
  • Word co-occurrence patterns within the corpus can therefore be explained through mixtures of such latent topics.
Formally, the likelihood of BTM is determined by the topic distribution and the topic–word distributions. Two sets of parameters must therefore be estimated: the topic proportion vector $\bm{\theta}$ and the topic–word distributions $\bm{\phi}_{z}$. The prior for each $\bm{\phi}_{z}$ is specified as a Dirichlet distribution with hyperparameter $\beta$, while the prior for $\bm{\theta}$ follows a Dirichlet distribution with hyperparameter $\alpha$. A latent topic assignment variable $z$ is drawn from a Multinomial distribution with parameter $\bm{\theta}$, and conditional on $z$, each word is generated from $\bm{\phi}_{z}$. Hence, the parameters of interest are $\bm{\theta}$, $\bm{\phi}_{z}$, and $z$.
The generative process of BTM can be summarized as:
  1. Step 1
    Draw a topic distribution $\bm{\theta}\sim\text{Dirichlet}(\alpha)$.
  2. Step 2
    For each biterm $b\in B$, assign a latent topic $z\sim\text{Multinomial}(\bm{\theta})$.
  3. Step 3
    For each topic $z$, draw a topic–word distribution $\bm{\phi}_{z}\sim\text{Dirichlet}(\beta)$.
  4. Step 4
    Generate the two words $w_{i},w_{j}\sim\text{Multinomial}(\bm{\phi}_{z})$.
The resulting joint likelihood over all biterms $B$ is:
$p(B)=\prod_{i,j}\sum_{z}\bm{\theta}_{z}\,\bm{\phi}_{i|z}\,\bm{\phi}_{j|z}.$ (1)
The conditional posterior for topic assignment $z$ is:
$p(z|\mathbf{z}_{-b},\mathbf{B},\alpha,\beta)\propto(n_{z}+\alpha)\frac{(n_{w_{i}|z}+\beta)(n_{w_{j}|z}+\beta)}{(\sum_{w}{n_{w|z}}+M\beta)^{2}},$ (2)
where $n_{z}$ is the number of biterms currently assigned to topic $z$, $n_{w|z}$ is the number of times word $w$ is assigned to topic $z$, and $\mathbf{z}_{-b}$ denotes topic assignments excluding the current biterm. To address convergence issues that may arise from direct Gibbs sampling, we employ collapsed Gibbs sampling, which integrates out $\bm{\theta}$ and $\bm{\phi}_{z}$ by exploiting conjugacy of the Dirichlet–Multinomial distributions [13]. After sufficient sampling iterations, the posterior estimates of topic–word distributions and topic proportions are:
$\bm{\phi}_{w|z}=\frac{n_{w|z}+\beta}{\sum_{w}{n_{w|z}}+M\beta},\quad\bm{\theta}_{z}=\frac{n_{z}+\alpha}{|B|+K\alpha}.$ (3)
Algorithm 1 Collapsed Gibbs Sampler for BTM
Input: number of topics $K$, hyperparameters $\alpha$, $\beta$, biterm set $\mathbf{B}$
Output: topic–word distributions $\bm{\phi}_{w|z}$ and topic proportions $\bm{\theta}_{z}$
1:Initialize topic assignments randomly for all biterms
2:for iteration $1,2,\ldots,N$ do
3:  for each biterm $b\in\mathbf{B}$ do
4:   Sample $z_{b}$ from $p(z|\mathbf{z}_{-b},B,\alpha,\beta)$
5:   Update counts $n_{w|z}$ and $n_{z}$
6:   Compute parameters $\bm{\phi}_{w|z}$ and $\bm{\theta}_{z}$
7:  end for
8:end for
After inference, the BTM yields both $\bm{\phi}_{w|z}$ and $\bm{\theta}_{z}$. Each topic is defined by its associated words and their probabilities, making it possible to compare topics by identifying their distinguishing word distributions. Because many words may appear across multiple topics with non-negligible probabilities, we focus on representative terms. Specifically, we construct a word–topic probability matrix $\mathbf{X}$ of dimension $N\times P$, where $N$ is the number of words and $P$ is the number of topics. To identify characteristic words, we compute two measures for each row of $\mathbf{X}$: (i) the coefficient of variation across topics, and (ii) the maximum probability. The coefficient of variation captures the relative dispersion of a word’s probabilities across topics, while the maximum probability identifies whether the word is strongly associated with at least one topic. Words with both high dispersion and high maximum probability are retained as representative terms, enabling more interpretable characterization of latent topics.
2.3 Clustering Topics via Latent Space Item Response Model
We estimate interactions among topics and visualize their relationships by embedding them into a latent interaction map. To achieve this, we follow the approach of Jeon et al. [9], who applied LSIRM to topic–word distributions, and employ its Gaussian version in our setting. Specifically, we use the Gaussian LSIRM to represent the bipartite structure between topics and words, where topics are regarded as “items” and words as “respondents” as below:
[…]
where $x_{i,p}$ denotes the probability of word $i$ belonging to topic $p$, $\bm{a}_{i}$ and $\bm{b}_{p}$ are attribute parameters, and ${\bf v}_{i}-{\bf u}_{p}||$ captures the Euclidean distance between the latent positions of topic $p$ and word $i$. A shorter distance implies stronger association between the word and the topic. Bayesian inference is employed to estimate the full parameter set $\bm{\Theta}=\{\bm{a},\bm{b},\mathbf{U},\mathbf{V}\}$ with appropriate priors, and parameters are sampled using Markov chain Monte Carlo (MCMC). As a result, we obtain latent positions of topics $\mathbf{v}_{i}$ in $\mathbb{R}^{d}$, forming the topic coordinate matrix ${\bf A}\in\mathbb{R}^{d\times P}$.
Once the latent positions of topics are estimated, we proceed to cluster the topics in order to identify groups with similar semantic characteristics. To this end, we apply the K-means clustering algorithm to the latent position matrix $\mathbf{A}$. The K-means method partitions the $P$ topics into $C$ disjoint clusters $\{C_{1},\ldots,C_{C}\}$ by minimizing the within-cluster sum of squared distances:
$\underset{C_{1},\ldots,C_{C}}{\arg\min}\sum_{c=1}^{C}\sum_{\mathbf{v}_{i}\in C_{c}}\left\lVert\mathbf{v}_{i}-\bm{\mu}_{c}\right\rVert^{2},$ (5)
where $\bm{\mu}_{c}$ denotes the centroid of cluster $C_{c}$. This clustering step groups topics that are located close together in the latent space, thereby reflecting their semantic similarity as derived from the topic–word distributions.
The resulting framework enables us not only to visualize relationships among topics through their latent embeddings, but also to categorize them into interpretable clusters. Specifically, LSIRM provides a probabilistic mechanism to embed topics in a common latent space, and K-means clustering on the estimated positions further organizes these topics into coherent groups. This joint approach allows us to explore both the global structure (via the interaction map) and the local grouping (via cluster assignments) of topics inferred from the text data.
2.4 Risk Index Factor via Lasso Regression
To assess the contribution of words in explaining fire-related financial losses, we model the expected property damage amount $y$ for each incident report as the outcome variable and the extracted words as predictors $x$. Specifically, let $\mathbf{y}=(y_{1},\ldots,y_{n})$ denote the vector of estimated damages across $n$ reports, and let $\mathbf{X}\in\mathbb{R}^{n\times p}$ be the document–term matrix, where each column corresponds to one of the $p$ words. We fit a Lasso regression model [19] of the form
$\hat{\bm{\beta}}=\arg\min_{\bm{\beta}}\Bigg\{\frac{1}{2n}\lVert\mathbf{y}-\mathbf{X}\bm{\beta}\rVert_{2}^{2}+\lambda\lVert\bm{\beta}\rVert_{1}\Bigg\},$ (6)
where $\lambda>0$ is a tuning parameter. The $\ell_{1}$ penalty induces sparsity in the estimated coefficients, effectively performing variable selection by shrinking many coefficients toward zero. This property enables us to identify only those words that exhibit substantial predictive power for property damage amounts.
The estimated coefficients $\hat{\beta}_{j}$ quantify the influence of word $j$ on expected fire-related losses. Coefficients close to zero indicate little or no association between the word and the damage amount, while coefficients with larger magnitude capture stronger effects. Moreover, the sign of $\hat{\beta}_{j}$ allows direct interpretation: negative values imply that the presence of a word is associated with lower property damage, whereas positive values suggest that the word signals higher expected losses. In this way, the Lasso framework provides a principled mechanism to evaluate and interpret the importance of words in the context of fire incident reports. Building on the estimated coefficients, we further construct a Risk Index to quantify word-level contributions in a structured manner. Specifically, we derive the index along three complementary dimensions.
First, within each cluster of semantically related words, the estimated coefficients are rescaled using a min–max transformation to lie between 0 and 1. This yields the i) Risk Index within Cluster ($\bm{\gamma}_{i,c},i=1,\cdots,n_{c},c=1,\cdots,C$), which allows for comparison of words relative to others in the same group. Second, to assess risk at the cluster level, we compute the mean coefficient value across words in each cluster $c$ and again apply min–max scaling to obtain a ii) Cluster-Level Risk Index between 0 and 1 ($\bm{\delta}_{c},c=1,\cdots,C)$. This provides an interpretable measure of the overall risk associated with each cluster. Finally, the iii) Overall Risk Index ($\bm{\rho}_{i,c}$) for a word is defined as the average of its within-cluster score and the risk score of its corresponding cluster. In this way, the framework jointly accounts for both the local importance of a word relative to its peers and the broader risk tendency of its semantic group. The resulting index serves as a principled indicator to identify key linguistic factors associated with higher fire-related property damages in incident reports.
3 Results
3.1 Topic Characterization through Words
Table 1 summarizes the thematic interpretation of topics extracted from the fire investigation records. In addition, Table 2 presents the distribution of words, listing the top five terms with the highest probabilities for each topic. These high-probability keywords serve as representative indicators of topic relevance and provide the basis for characterizing the underlying thematic structure.
Table 1: Mapping of extracted topics from fire investigation records to their descriptive names. The thematic categories are assigned based on high-probability keywords (summarized in Table 2) and domain-specific interpretation.
Topic Name Topic 1 Flammable vapor ignition due to the use of organic solvents such as cleaning solutions Topic 2 Explosions caused by sparks generated from friction or static electricity due to the blending of flammable or combustible raw materials Topic 3 Fires caused by common electrical factors Topic 4 Ignition caused by sludge accumulation within ventilation equipment Topic 5 Ignition caused by electrical heating Topic 6 Related to dust collection equipment Topic 7 Spontaneous combustion caused by improper storage or containment of flammable residues Topic 8 Fire caused by forklift operation Topic 9 Chemical explosion occurring in the reactor Topic 10 Ignition of residue accumulated in ducts connected to or adjacent to dust collection equipment Topic 11 Due to improper use of drying equipment Topic 12 Fire caused by improper use of a banbury mixer for rubber molding Topic 13 Combustible materials ignited due to welding spark during hot work Topic 14 Fire caused by ignition sources in machinery with operation motor, such as air compressors Topic 15 Ignition of waste materials due to improper disposal of cigarette butts
Topic 1 reflects incidents associated with oil vapors generated from organic solvents or related chemical processes, frequently occurring in cleaning operations or wastewater treatment facilities. Topic 2 captures fire risks arising from the ignition of combustible materials due to friction or static electricity during equipment operation, as suggested by keywords such as mixer and drum can. Topic 3 represents general electrical fires, characterized by terms such as electrical short circuit and circuit breaker, while Topic 5 is more narrowly related to electrical heating sources (e.g., air conditioners and heating wires). Topic 4 emphasizes ignition triggered by sludge accumulation within ventilation systems, particularly in laboratory environments.
Topic 6 highlights fires directly linked to dust collection equipment, where flames may propagate through ducts or filter systems containing combustible dust. In contrast, Topic 10, though conceptually related, places less emphasis on dust collection equipment itself and instead indicates fire hazards involving adjacent facilities, such as ventilation ducts, plastics, and drying equipment. Topic 7 is dominated by spontaneous combustion events caused by the improper storage of combustible residues, including processed byproducts such as sesame dregs. Topic 8 involves forklift-related fires, which occur across both chemical and general factory settings, often due to battery or engine compartment failures. Topic 9 concerns reactor-related incidents in chemical plants, where abnormal reactions generate oil vapors leading to explosions.
Other topics describe equipment-specific or context-specific fire causes. Topic 11 focuses on drying equipment, particularly in cases involving powders such as silicon. Topic 12 reflects fires ignited by the thermal oil of banbury mixing equipment. Topics 13 through 15 capture more general factory-related accidents: Topic 13 refers to welding-induced ignition near cooling towers or sandwich panels; Topic 14 highlights motor-related electrical fires in air compressors; and Topic 15 illustrates landfill or waste-area fires, frequently initiated by discarded cigarette butts.
While biterm modeling provides an effective means of partitioning documents into topics, it assigns every word to all topics, which limits its ability to capture direct relationships among words. Since the primary goal of this study is to identify the words that carry substantial meaning within fire investigation documents, it is essential to explore how words are grouped through their interactions. In this respect, topic–word distributions offer probabilistic information on the degree of association between words and topics, thereby serving as a valuable resource for indirectly inferring inter-word relationships. Leveraging this information enables us to examine documents not only at the level of topics, but also from the perspective of word-level associations, ultimately facilitating a more interpretable summarization of fire-related narratives.
Table 2: Top five representative words for each extracted topic based on topic–word probability distributions. These high-probability words provide the basis for interpreting the thematic characteristics of the topics.
Topic Top 5 Words Topic 1 cleaning, cleanig room, wastewater treatment plant, agitator, machine room Topic 2 mixer, drum can, pallet, plastic, explosion Topic 3 electrical short-circuit, electrical distribution, flame, circuit breaker, SWGR Topic 4 extractor hood, oven, laboratory, transformer, electricity Topic 5 heating wire, outdoor unit of air-conditioner, thermal/acoustical insulation, air-conditioner, rooftop Topic 6 dust collection equipment, duct, flame, plenty of, filter system Topic 7 storage, sesame dregs, residues, spontaneous combustion, storage room Topic 8 forklift, battery, electrical short-circuit, distribution, engine compartment Topic 9 explosion, container, reactor, oil vapor, static electricity Topic 10 dust collection equipment, duct, flame, ventilation duct, plastic Topic 11 drying equipment, powder, silicon, base material, storage Topic 12 base material, mixing equipment, banbury, thermal oil, flame Topic 13 welding, cooling tower, flame, sandwich panel, acetylene Topic 14 air compressor, motor, electrical short-circuit, vulcanizer, solenoid valve Topic 15 waste materials, cigarette butts, waste, stacking, flame
3.2 Thematic Aggregation of Topics
Building upon the topic–word distributions, we further inferred interactions among words to explore higher-level thematic structures. Figure 2 illustrates the latent positions of words projected onto a two-dimensional space, where clustering was performed using the $k$-means algorithm. The visualization highlights clusters by distinct colors, showing how semantically related words are grouped in close proximity. Based on model selection criteria, a total of 15 clusters were identified as valid. As shown in the figure, words located near one another in the latent space tend to form coherent clusters, thereby capturing meaningful associations beyond the topic-level representation.
Figure 2: Two-dimensional latent representation of words with clustering results obtained using the $k$-means algorithm. Distinct colors indicate the 15 identified clusters, showing how semantically related words are grouped in close proximity.
The interpretation of clusters requires careful consideration because the meaning of a single word is better understood in relation to its neighboring terms. A word located at the center of a cluster typically exhibits a high probability of contributing to the generation of a topic, often alongside other words in the same group. However, the semantic similarity within a cluster may arise either from the joint contribution of the words themselves or from their shared functional context with alternative terms. Thus, rather than examining individual words in isolation, it is crucial to evaluate their surrounding vocabulary to delineate the collective meaning of each cluster. In this sense, clusters serve as categorical units in which semantic coherence emerges from localized word proximities, allowing for the clarification of latent thematic structures.
3.3 Cluster Interpretation
According to Table 3, the top 10 keywords in each cluster are ordered by their proximity to other words, as determined by distance calculations based on the estimated latent positions. This ordering enables a more precise characterization of the distinctive features of each cluster. The identified clusters can be grouped into several overarching thematic categories of fire risks.
Chemical and Reaction-Related Risks.
Clusters involving hazardous chemical processes highlight the potential for fires and explosions arising from chemical leaks, abnormal reactions, or spontaneous combustion of unstable substances. For instance, distillation column, exposure, and solvent (Cluster 1) point to explosions due to hazardous material leakage and pressure control failures. Similarly, abnormal reaction terms such as pharmaceutical, film, and epoxy (Cluster 5) suggest severe accidents during abnormal chemical handling. Risks of spontaneous combustion are also evident, as indicated by heat of reaction, expired reagents, and cooking oil (Cluster 4), as well as residues like compost and sesame dregs (Cluster 14), which underscore dangers from abandoned chemicals and oils.
Industrial Work and Machinery-Related Risks.
Another group of clusters reflects ignition sources associated with industrial work environments and machinery. Terms such as polishing, painting, and waste (Cluster 2) emphasize ignition from heat accumulation near combustible interior materials. The presence of incinerate, waste wood, and dust collection equipment (Clusters 3 and 15) indicates residual heat or sparks generated during work with combustibles. Likewise, coating, cutting oil, and grinders (Cluster 7) highlight ignition risks from friction heat or sparks produced by machinery.
Electrical and Equipment-Related Risks.
Electrical fire hazards are represented by clusters centered on terms such as storage, cable, and control box (Cluster 6), which indicate electrical fires in storage or handling areas. Other clusters highlight process-related electrical causes (Cluster 11) and fires occurring during maintenance of machinery (Cluster 12). These categories capture risks arising not from chemical reactions but from equipment and power systems in industrial sites.
Vapor, High-Temperature, and Dust Accumulation Risks.
A further set of clusters concerns ignition in environments with flammable vapors, high temperatures, or combustible dust. For example, vapor, hazardous materials, and grease (Cluster 8) illustrate ignition hazards during the use of organic solvents. Likewise, drying room, cosmetic, and ventilation fan (Cluster 9) point to fire risks in areas with rising oil vapor temperatures. Clusters including dilution, melting fusion, and vulcanizer (Cluster 13) capture ignition due to accumulated combustible dust and oil vapors, further extending the scope of hazards beyond purely chemical origins.
Accumulated Combustibles.
Finally, some risks are associated with the buildup of general combustible materials, as indicated by Cluster 10, which represents fire outbreaks triggered by ignition in accumulated combustibles. This category is particularly relevant to non-chemical industrial environments where storage and waste management play a central role.
Table 3: Clusters of 10 representative keywords derived from inter-word correlations based on estimated latent positions.
Cluster 10 Words 1 distillation column, exposure, solvent, photoinitiator, metal, cleaning solvent, heptane, nozzle, flexible, silanes manufacturing 2 ignition source, polishing, painting, floor, laboratory, electricity, plenty of, scrap paper, waste, recycling 3 incinerate, waste wood, interior material, heat, urethane, smoldering ignition, cutting machine, dust collection equipment, rubber, manufacturing machine 4 heat of reaction, expired reagent, corn, cooking oil, storage tank, reagent, oxygen, eruption, heat wave, rainwater 5 abnormal reaction, pharmaceutical, film, sheath heater, upper, suction, oil, wire mesh, epoxy, dust explosion 6 storage, cable, malfunction, cool down, oil tank, control box, coil, electrical circuit board, refrigerator, fertilizer 7 coating, injection, corrosion, cutting oil, grinder, press, repair, acetylene, scrubber, steam equipment 8 vapor, nucleic acid, grease, leak, hazardous material, mix, paint, high temperature, large-scale, thermal cutting 9 drying room, cosmetic, pallet, ventilation fan, pressure, ceiling, rotation, small amount, decompose, lid 10 preheat, ignition, stacked, accumulation, tire, moisture, activated carbon, semi-finished product, waste storage, welding 11 cooling fan, service line, shut out, lower terminal, insulating oil, underground, fan belt, switch, electric current, analysis lab 12 closing, charging equipment, flammability, maintenance, automation, repair, cooling machine, machine room, gas torch, arc 13 dilution, power outage, cover, immediate upper, tracking, prefabricated, vacuum, muller, vulcanizer, melting fusion 14 micro, seat pad, deodorizing tower, absorbent pad, gunnysack, ton bag, compost, boiler room, sesame dregs, oxidation heat 15 belt, particle, capture, aluminum, inlet, rubber department, wood chip, blowing, metal powder, abradant
Taken together, the interpretation of clusters provides meaningful categorical insights into the semantic structure of fire investigation records. However, semantic similarity alone does not capture the extent of economic severity associated with each term. To address this, we incorporated fire-damage estimates into the analysis by estimating word-level coefficients via LASSO regression. This approach allows us to link clusters with the magnitude of potential financial losses. The coefficients estimated for each word quantify its relative contribution to explaining variations in property damage estimates. Figure 3 presents a three-dimensional visualization, where the horizontal axes represent the latent positions of words and the vertical axis corresponds to their regression coefficients. The clusters are indicated by color, consistent with the grouping shown in Figure 2.
As illustrated, even within the same semantic cluster, words exhibit heterogeneous patterns: some words have positive coefficients, indicating stronger associations with larger property losses, while others display negative coefficients, reflecting lower associated damage levels. This observation highlights that clusters capture thematic similarity but do not necessarily imply uniform economic consequences. Hence, incorporating these regression coefficients into subsequent analyses provides an additional, economically grounded perspective. In particular, the integration of semantic clustering with property damage-based coefficients motivates the construction of a risk index that simultaneously reflects linguistic structure and financial severity, thereby offering a more comprehensive measure of fire-related risks.
Figure 3: Three-dimensional visualization of words, where the $x$-and $y$-axes represent latent positions and the $z$-axis corresponds to coefficients estmated from a LASSO regression on fire-damage insurance claims. Colors indicate clusters identified through $k$-means grouping.
3.4 Risk Index Estimation
The preceding analyses demonstrate that clusters derived from topic–word distributions provide semantically coherent categories of fire-related terms, while regression coefficients estimated from property damage amounts capture the associated economic severity. Importantly, these two perspectives highlight complementary aspects of risk: semantic clusters reflect the contextual mechanisms of fire occurrence, whereas coefficients quantify their financial impact. Moreover, as shown in the regression analysis, even words within the same cluster may exhibit heterogeneous patterns of association with damage amounts, indicating that semantic similarity alone is insufficient for fully characterizing fire risk.
To address this limitation, we propose the construction of a composite risk index that integrates both linguistic and economic dimensions. To quantify the relative contribution of words and clusters to fire-related incidents, we define three levels of risk indices: the word-level index $\gamma_{i,c}$, the cluster-level index $\delta_{c}$, and the overall word index $\rho_{i,c}$. Each measure captures a distinct dimension of risk, ranging from fine-grained lexical associations to broader thematic categories. By jointly considering (i) cluster-level semantic associations and (ii) word-level coefficients derived from loss data, the risk index provides a systematic measure of fire risk that is interpretable in terms of language use and grounded in economic outcomes.
Word-level Risk Index ($\gamma_{i,c}$).
The index $\gamma_{i,c}$ measures the relative contribution of word $i$ within cluster $c$ to the estimation of property damage amounts. A higher value indicates that the word is more strongly associated with larger expected losses compared to other words in the same cluster. Table 4 presents the top ten words with the highest $\gamma_{i,c}$ values in each cluster, along with their Risk Index. For example, in Cluster 2 (Ignition from heat accumulation near combustible interior materials) and Cluster 7 (Ignition of flammable substances by friction heat or spark from machinery), words such as flame, heat, and grinder show high $\gamma_{i,c}$ values, reflecting their strong linkage with higher levels of estimated property damage. Thus, $\gamma_{i,c}$ highlights words whose relative importance provides insight into the financial risk implications captured within each cluster.
Cluster Word 1 Word 2 Word 3 Word 4 Word 5 Word 6 Word 7 Word 8 Word 9 Word 10 1 photoinitiator (1.000) ethyl acetate (0.911) reactor (0.846) heptane (0.814) sclerotic (0.809) centrifuge (0.798) lesk (0.781) cleaning solvent (0.741) solvent (0.739) chemical reaction (0.736) 2 flame (1.000) plenty of (0.937) wall (0.893) interior wall (0.891) laboratory (0.858) floor (0.856) radiant heat (0.834) laboratory (0.736) insulator (0.712) exterior wall (0.674) 3 dust collection equipment (1.000) urethane (0.907) smoldering ignition (0.740) rubber (0.714) interior material (0.657) heat (0.493) cutting machine (0.445) waste wood (0.082) incinerate (0.000) – (1.000) 4 stacked goods (0.987) oil vapor (0.970) storage tank (0.675) gas (0.599) oxygen (0.581) reagent (0.448) dust collector (0.427) waste (0.381) dummy (0.317) oil (0.296) 5 manhole (1.000) expander (0.949) cover (0.939) film (0.890) epoxy (0.819) mixer (0.801) manufacturing building (0.798) static electricity (0.683) impurities (0.633) physic (0.594) 6 shipping area (1.000) petroleum product (0.879) condenser (0.878) storage (0.745) cable (0.708) refrigerator (0.691) arson (0.642) crayon (0.539) oil tank (0.484) hazardous substances plant (0.461) 7 grinder (1.000) polystyrene (0.952) fabric (0.909) long time (0.902) coating (0.816) demolition (0.780) manufacturing room (0.769) panel (0.747) electrical distribution (0.648) cutting oil (0.637) 8 chemical material (1.000) hazardous material (0.852) leak (0.753) vapor (0.680) metal plate (0.603) base material (0.544) nucleic acid (0.456) spontaneous combustion (0.424) grease (0.402) large-scale (0.341) 9 warehouse (1.000) pallet (0.933) storage (0.851) cosmetic (0.825) agitator (0.789) drying room (0.712) liquid (0.696) secondary battery (0.693) staff lounge (0.665) heater rod (0.522) 10 activated carbon (1.000) welding (0.863) remain (0.816) moisture (0.588) semi-finished product (0.572) accumulation (0.395) tire (0.393) paint (0.378) plastic (0.369) ignition (0.367) 11 service line (1.000) packing room (0.869) analysis lab (0.841) interlayer short circuit (0.833) distribution board (0.826) molding machine (0.712) unidentified short circuit (0.673) power line (0.656) trip (0.618) air-compressor (0.616) 12 automation (1.000) flammable (0.938) repair (0.843) machine room (0.607) insulation deterioration (0.541) air-conditioner (0.356) arc (0.328) terminal (0.316) battery (0.268) outdoor unit of air-conditioner (0.258) 13 vulcanizer (1.000) prefabricated (0.662) research building (0.553) indoor wiring (0.500) dilution (0.492) complete product (0.482) automobile parts (0.480) filtering equipment (0.465) boiler (0.460) circuit breaker (0.455) 14 roof (1.000) entrance (0.969) unidentified cause (0.889) machine (0.885) sandwich panel (0.829) appliances (0.768) dormitory (0.746) compost (0.734) boiler room (0.634) deodorizing tower (0.609) 15 lower (1.000) explosion (0.769) electrical equipment (0.684) commissioning (0.622) duct (0.621) rigid polyurethane (0.605) stacked materials (0.595) melting equipment (0.574) condensation (0.562) freeze drier (0.548)
Table 4: Cluster–Word table with Risk Index values. For each cluster, the top ten words with the highest word-level Risk Index ($\gamma_{i,c}$) are reported. These words represent the relatively high-scoring lexical elements within each cluster, indicating stronger contributions to the estimated property damage amounts.
Cluster-level Risk Index ($\delta_{c}$).
The index $\delta_{c}$ summarizes the risk associated with cluster $c$ as a whole, aggregating the estimated coefficients of its constituent words. A higher $\delta_{c}$ indicates that, on average, words belonging to this cluster are strongly predictive of higher property losses. Table 5 reports these values. The five clusters with the highest $\delta_{c}$ include: (i) Ignition of flammable vapor during organic solvent use, (ii) Electrical fires in areas handling flammable materials, (iii) Ignition from temperature rise in oil vapor areas, (iv) Ignition from heat accumulation near combustible interior materials, and (v) Fires or explosions from abnormal reactions during chemical handling. These topics correspond to scenarios where ignition sources and flammable environments directly translate into severe financial consequences. By contrast, the clusters with the lowest $\delta_{c}$, such as Ignition from sparks inside dust collection equipment with filters and debris or Fires from electrical factors in process equipment sites, represent situations with relatively weaker association to large-scale losses. In this way, $\delta_{c}$ provides an interpretable measure of how strongly each cluster of fire-related factors contributes to financial risk.
Cluster Risk Index Topic
8 1.000 Ignition of flammable vapor during organic solvent use
6 0.954 Electrical fires in areas handling flammable materials
9 0.731 Ignition from temperature rise in oil vapor areas
2 0.639 Ignition from heat accumulation near combustible interior materials
5 0.626 Fires or explosions from abnormal reactions during chemical handling
12 0.606 Fires during machinery maintenance
7 0.561 Ignition of flammable substances by friction heat or spark from machinery
4 0.534 Spontaneous combustion from abandoned chemicals and oils
10 0.405 Fire from ignition in accumulated combustibles
13 0.388 Ignition in areas with accumulated combustible dust and oil vapors
14 0.344 Spontaneous combustion from improper oil residues storage or disposal
3 0.294 Ignition due to residual heat post-work with combustibles
1 0.216 Fire or explosion caused by chemical leakage
15 0.205 Ignition from sparks inside dust collection equipment with filters and debris
11 0.000 Fires from electrical factors in process equipment sites
Table 5: Cluster ranking by Risk Index values and their associated topics. The Risk Index ($\delta_{c}$) represents the aggregated risk contribution of each cluster, with higher values indicating strong fire-related financial losses.
Overall Word Risk Index ($\rho_{i,c}$).
The index $\rho_{i,c}$ extends beyond cluster membership to evaluate the global risk contribution of word $i$, accounting for its position across the latent embedding space. Table 6 lists the top twenty words by $\rho_{i,c}$. For example, chemical material, shipping area, and hazardous material emerge as the top three words. These terms directly connect to concrete accident scenarios such as the generation of flammable vapors from sludge leakage, electrical fires in distribution boards, and vapor leakage during hazardous material processing. The $\rho_{i,c}$ index therefore highlights words that carry not only lexical salience within clusters but also broader cross-cluster risk relevance.
Collectively, these three indices provide a multi-layered framework: $\gamma_{i,c}$ identifies salient words within clusters, $\delta_{c}$ ranks the clusters by their aggregate hazard potential, and $\rho_{i,c}$ detects globally critical words linked to real-world accident narratives.
No. Word Risk Factor 1 chemical material 1.000 2 shipping area 0.977 3 hazardous material 0.926 4 petroleum product 0.917 5 condenser 0.916 6 warehouse 0.866 7 storage facility 0.850 8 pallet 0.832 9 cable 0.831 10 flame 0.820 11 manhole 0.813 12 automation 0.803 13 storage 0.791 14 plenty of 0.788 15 expander 0.788 16 cosmetic 0.778 17 stacked goods 0.761 18 agitator 0.760 19 film 0.758 20 spontaneous combustion 0.712
Table 6: Word importance based on overall Risk Index ($\rho_{i,c}$). The table reports the top 20 words with the highest $\rho_{i,c}$ values, highlighting globally influential keywords strongly linked to insurance loss outcomes.
4 Discussion
Wehmeier and Mitropetrosb [22] analyzed the causes of fires in chemical plants based on incidents at a chemical–pharmaceutical company and categorized them as follows: Self-ignition (22%), Hot running of moving parts (17%), Welding (15%), Electrostatic (14%), Drying (10%), Repair/Maintenance (8%), Leakage (7%), and Electric (short-circuit) (7%). However, he also concluded that fire investigations in the German chemical industry reveal complex and heterogeneous causal structures. Furthermore, because this analysis relied solely on accident-frequency statistics, it was insufficient for assessing risk in the sense of loss severity. Complementarily, Darbra et al. [4] constructed a relative-probability event tree for major chemical accidents with domino effects, analyzing the causes, materials involved, effects and consequences, affected population, and the likelihood of specific accident sequences, thereby emphasizing the need to evaluate interconnected hazards rather than isolated causes.
To address these limitations, we replace frequency-based tallies with a composite, loss-aware scoring framework that links text-derived indicators to the scale of damage. The procedure is twofold: (i) estimate a risk score for each term that reflects its association with property-relevant losses, and (ii) interpret high-scoring terms by examining their local semantic neighborhoods to recover operational and environmental context.
We instantiate this framework via the Overall Word Risk Index, $\rho_{i,c}$, which quantifies the association between terms and loss outcomes rather than raw frequency. Using $\rho_{i,c}$ to rank and organize the vocabulary in the embedding (Fig. 4), we identify four cross-cutting facets of fire risk: (i) chemical leakage/vapors (red), (ii) storage/warehouses (blue), (iii) equipment/electrical (green), and (iv) self-ignition (pink). In what follows, we analyze each facet and its representative cases.
Four Cross-cutting Facets of Fire Risk
Figure 4: Two-dimensional embedding (v1, v2) of fire-risk keywords, colored by four discussion facets and shaped by cluster assignments. Colored, labeled points are representative high-scoring terms (ranked by the Overall Word Risk Index $\rho_{i,c}$) in parentheses. Marker shapes encode cluster names as shown in the legend; gray points denote the remaining vocabulary.
First, the highest-ranking terms such as chemical material, hazardous material, and petroleum product capture risks associated with chemical leakage and the generation of flammable vapors. These findings align with well-documented hazards in chemical industries, where improper handling, leakage, or inadequate containment of chemical substances often lead to large-scale fire and explosion events. The 2009 Jaipur crude oil pipeline tank fire in India is an example of how a fire originating from a hazardous material leak can spread. The accident resulted in 13 deaths and over 200 injuries, setting a record for the worst accident of its kind in India [12]. The representative cases illustrate how unintended chemical leaks can escalate into severe incidents with significant property and human losses.
Second, structural and operational factors such as warehouses, storage facilities, and pallets are also prominent, underscoring the critical role of storage conditions and facility maintenance. Prior research on cold-chain logistics demonstrates that pallet stacking and package design strongly influence airflow and heat transfer within storage rooms, where insufficient ventilation can generate localized hotspots and elevate fire hazards [1]. In the context of energy storage, investigations highlight that inadequate cooling and airtight designs in portable energy storage systems promote overheating and exacerbate ignition risk [6]. Similarly, studies on lithium-ion batteries reveal that poor thermal management and accumulation of flammable gases in confined storage spaces can accelerate thermal runaway and combustion [7]. Together, these findings indicate that storage practices—ranging from palletized materials to battery warehouses—interact with flammable substances in ways that compound ignition potential, while recurring references to ventilation-related equipment emphasize how failures in airflow management can amplify both the likelihood and severity of fire incidents.
Third, a notable set of keywords points to ignition sources related to malfunctions or inadequate maintenance of electrical and mechanical equipment, such as condenser, cable, automation. These emphasize that not only causes directly related to chemical processes, but also the integrity of auxiliary mechanical and electrical systems, play a crucial role in fire risk. Electrical fires are among the most universal causes of fire. Korean statistics show that 27.4% of fires incidents from 1996 to 2021 were attributed to electrical factors [11], suggesting that the chemical industry is no exception to this risk. These findings suggest that the damage resulting from contact between common ignition sources and flammable raw materials or products handled in chemical plants can be significantly greater. The cases show that malfunctioning condensers, short-circuited cables, or maintenance using gas torches can trigger fires even in otherwise controlled environments.
Finally, terms such as spontaneous combustion and stacked goods reflect risks arising from self-ignition processes of chemicals and oil residues produced or derived in chemical plants, as well as environmental conditions. Recent studies on biomass storage have shown that moisture exchange, oxygen penetration, and heat accumulation can interact to induce self-heating and ultimately spontaneous ignition [23]. Similarly, investigations in mine waste dumps highlight that spontaneous combustion can compromise geomechanical stability, illustrating how natural or reactive processes extend the scope of fire hazards beyond purely mechanical or chemical failures [15]. These findings underscore the importance of incorporating self-ignition phenomena into fire risk assessments for chemical facilities and storage environments.
Taken together, these findings demonstrate that high-risk keywords are not confined to a single category but span across chemical substances, storage practices, equipment reliability, and environmental interactions. By transforming unstructured fire investigation narratives into analyzable data, our framework systematically extracted semantically related terms through latent topic and embedding models, contextualized them with representative fire cases, and linked them to property damage estimates. This integrative approach culminated in the development of a composite Risk Index, which quantifies the relative importance of text-derived indicators in relation to financial loss outcomes. The combination of statistical evidence with real-world loss information provides interpretability and practical relevance, showing how latent textual patterns can be operationalized into measurable safety metrics. In doing so, the proposed risk index offers a model-based tool for identifying actionable risk factors that bridge large-scale textual evidence with practical fire safety management.
Risk Index Perspective
As summarized in Table 4, the chemical leakage/vapors facet concentrates the highest $\rho_{i,c}$ values (e.g., chemical material, hazardous material, petroleum product), indicating a stronger linkage to larger financial losses in our corpus—consistent with evidence that leakage and vapor formation are prominent drivers of severe incidents in chemical industries [22, 12]. By contrast, the self-ignition facet (e.g., spontaneous combustion, stacked goods) exhibits lower $\rho_{i,c}$ values relative to the chemical-leakage/vapors facet in our corpus. This attenuation is consistent with three data-plausible mechanisms supported by prior work. First, self-heating often proceeds as a long-duration, low-temperature smouldering process with comparatively low heat-release rates and slow spread, which increases the opportunity for intervention before very large property losses accrue [16, 21, 18]. Second, consistent with prior surveys of major incidents, very large losses in hydrocarbon processing frequently originate from sustained leaks that evolve into flash fires or vapour cloud explosions, rather than from long-duration smouldering scenarios [2, 3, 4]. Third, contexts in which spontaneous combustion becomes catastrophic (e.g., biomass or mine-waste stockpiles) are emphasized in the engineering literature but are under-represented in our chemical special-building corpus, attenuating the empirical linkage between self-ignition terms and large losses [23, 15]. Interpreted this way, the relatively lower scores reflect dataset composition and event progression characteristics, rather than any contradiction with the established self-heating mechanism.
Why similarly clustered terms admit distinct operational readings.
Although cluster membership (shapes in Fig. 4) reflects lexical neighborhoods, nearby terms can encode different operational contexts that co-occur in narratives. For example, storage terms (blue) can sit beside equipment/electrical terms (green) because pallet stacking and enclosure geometry restrict ventilation and heat rejection [1], thereby increasing the efficacy of routine electrical faults as ignition sources; likewise, self-ignition terms (pink) may appear near leakage/vapor terms (red) when oxygen ingress and heat accumulation in stacked goods elevate vapor formation and ignition potential [23]. Conversely, process-centric terms (e.g., maintenance/automation) may be spatially offset when hot-work contexts (sparks, localized heating) are discussed apart from storage constraints, yet they remain semantically bridgeable whenever operations occur proximate to flammable inventories or confined airflow paths [5, 7]. Interpreting the embedding jointly with the loss-aware index $\rho_{i,c}$ (Fig. 4) discriminates semantically adjacent yet economically distinct patterns: terms that lie close in the map but carry higher $\rho_{i,c}$ mark contexts historically associated with larger property losses, whereas nearby low-$\rho_{i,c}$ terms indicate operational exposure without the same tail severity. This joint reading—proximity for mechanism, $\rho_{i,c}$ for consequence—yields a ranked set of actionable priorities, directing inspection toward leakage/vapor configurations while maintaining vigilance for storage and self-ignition scenarios that can escalate under adverse conditions.
Methodologically, our analysis proceeds in four stages. First, we transform unstructured investigation narratives into topical structure via topic modeling and latent embeddings, which provide low-dimensional coordinates for terms and documents. Second, we quantify word–word interactions by combining co-occurrence statistics with embedding-based proximity to form a network. Third, we partition this network into semantically coherent clusters and assign interpretable labels using representative cases to recover operational and environmental contexts. Fourth, we couple these text-derived signals with structured loss data variables to construct a loss-aware score—the Overall Word Risk Index $\rho_{i,c}$—that prioritizes terms by their association with financial outcomes. Interpreting the embedding jointly with $\rho_{i,c}$ thus yields a map that is both mechanistic (pathways reflected in neighborhoods) and consequential (loss-linked salience), enabling prevention, inspection, and maintenance efforts to focus on the most actionable risks.
Future research will extend the current framework in two directions. First, we plan to develop methods for systematically tracking the temporal trends of risk factors, enabling the identification of how the prevalence and severity of specific hazards evolve over time. Such analyses will provide an evidence-based foundation for monitoring emerging risks and for designing timely interventions. Second, we aim to advance causal inference approaches tailored to text-derived risk indicators, with a particular focus on risky keywords identified in investigation records. By establishing causal relationships rather than mere associations, this line of work will enhance the interpretability of the extracted factors and strengthen their utility for policy and decision-making in fire risk management.